# AI4Science: The Reach and Limits of AI for Scientific Discovery

### v1

@@ 1 Introduction
Molecular simulations are indispensable for studying the complex dynamics that govern biological function, yet brute-force approaches struggle to access the slow, functionally relevant motions—such as protein folding, ligand binding, or allosteric switching—due to rugged energy landscapes and the dominance of rare events (Dror et al., 2011; Latorraca et al., 2017; Wang et al., 2024; Zhang et al., 2024a). To mitigate this gap—beyond what enhanced sampling can offer when suitable collective variables are hard to specify—a complementary strategy has gained traction: shifting the burden from raw coordinates to learned latent coordinates. In this representation-first view, the simulation problem is recast as a modular encoder–propagator–decoder pipeline: an encoder maps high-dimensional atomic configurations into a continuous, low-dimensional latent space; a propagator evolves the system’s state within this simplified space; and a decoder maps the resulting latent trajectory back to all-atom coordinates (Sidky et al., 2020; Wu et al., 2018).
Progress within this paradigm comes from two complementary directions. The first focuses on learning the underlying physics, employing score-based diffusion, flow matching, and energy-based models to learn generative surrogates—sometimes yielding differentiable force fields—that implicitly define the system’s potential of mean force (Arts et al., 2023; Plainer et al., 2025). The second centers on learning simplified dynamical coordinates, using time-aware autoencoders or Koopman/DMD analysis to discover an intrinsic manifold where the long-term dynamics become stable, predictable, or even approximately linear (Hernández et al., 2018; Mardt et al., 2018; Brunton et al., 2021; Li et al., 2023a; Zhang et al., 2024b; Azencot et al., 2024).
@@ 2 Related Work
The core idea of simulating in a low-dimensional space is well established. Molecular Latent Space Simulators (LSS) (Sidky et al., 2020) explicitly factorize the problem into three components: an encoder to find slow collective variables (CVs), a latent propagator to evolve them, and a decoder to generate all-atom structures. Similarly, Deep Generative MSMs (DeepGenMSM) (Wu et al., 2018) pair a latent Markovian transition model with a generative decoder to emit molecular configurations for each state. This blueprint has been realized in various forms, including autoregressive simulators with RNN/LSTM propagators (Vlachas et al., 2021), trajectory-level generators that cast MD as video synthesis (Jing et al., 2024), and invertible models like Boltzmann Generators that learn a direct map to the equilibrium distribution (Noé et al., 2019). We follow this modular design but (i) hold the LD-FPG decoder fixed to control for reconstruction quality and (ii) focus our analysis on the choice of latent propagator (Fuchs et al., 2020; Satorras et al., 2021; Xu et al., 2022; Thölke and De Fabritiis, 2022).
@@
The quality of a latent simulation hinges on the quality of the latent space itself. While linear methods like time-lagged independent component analysis (TICA) remain strong baselines for identifying slow variables (Pérez-Hernández et al., 2013), deep learning has enabled far more expressive encoders. Time-lagged autoencoders such as the Variational Dynamics Encoder (VDE) learn non-linear representations predictive over a time delay $\Delta t$ (Hernández et al., 2018; Wehmeyer and Noé, 2018). VAMPnets use a variational principle to approximate the leading Koopman eigenfunctions, corresponding to the slowest dynamical processes (Mardt et al., 2018). To remove hand-crafted features, modern approaches leverage graph neural networks (GNNs) to learn CVs directly from coordinates in a permutation- and symmetry-aware manner (Zou et al., 2025; Ghorbani et al., 2022; Zhang et al., 2024c; Duval et al., 2023), often with information-bottleneck objectives to explicitly optimize predictiveness of future states (Wang et al., 2021).
@@
The Koopman-operator framework provides a route to linearize non-linear dynamics in a learned observable space (Brunton et al., 2021). Extended dynamic mode decomposition (EDMD) (Williams et al., 2015) and DMD (Tu et al., 2014) showed that a wide class of systems admits accurate linear predictors; subsequent work learns such observables end-to-end with deep encoders (Lusch et al., 2018). This has inspired Koopman autoencoders that enforce a linear evolution rule within the latent space (Tayal et al., 2023; Wu et al., 2025), yielding exceptional long-horizon stability by avoiding the compounding errors of iterated non-linear models, with continued advances in scalable kernels and consistency-enforcing architectures (Klus et al., 2023; Azencot et al., 2024).
@@
A complementary path directly learns the non-linear transition function. The Learning Effective Dynamics (LED) framework uses LSTMs to propagate latent variables and capture memory effects (Vlachas et al., 2021); others pair RNNs with physical resampling such as Maximum Caliber to enforce kinetic consistency (Tsai et al., 2022). Continuous-time learners (Neural ODEs) offer flexible parametric flows (Chen et al., 2018; Li et al., 2023b), and graph-based simulators exploit locality for rollouts in interacting systems (Sanchez-Gonzalez et al., 2020). To curb instability, physics-preserving architectures (Hamiltonian/Lagrangian networks) enforce conservation laws (Greydanus et al., 2019; Cranmer et al., 2020). Our MLP propagator serves as a streamlined baseline in this family.
@@
Generative models increasingly embed statistical-mechanics structure. Score-based diffusion has been used to learn effective force fields for coarse-grained MD (Arts et al., 2023), linking denoising, Langevin dynamics, and Fokker–Planck evolution (Song et al., 2021). Recent works (e.g., DiffMD, Score Dynamics) demonstrate larger stable time steps while retaining short-time kinetic signals (Wu and Li, 2023; Hsu et al., 2024). A key theme is consistency between the sampled equilibrium ensemble and the stationary distribution of the learned dynamics (Plainer et al., 2025); stability can also benefit from noise augmentation during training (Harcombe and Duignan, 2025). We adopt these ideas in a latent-space setting (De Bortoli et al., 2021; Raja et al., 2025).
@@
G protein-coupled receptors (GPCRs) are a demanding testbed: their function hinges on slow transitions among metastable states mapped by landmark MD/MSM studies (Dror et al., 2011; Shukla et al., 2014; Kohlhoff et al., 2014; Latorraca et al., 2017; Weis and Kobilka, 2018). High-resolution structures (e.g., $\beta_{2}$AR–Gs) provide anchors for validating pathways (Rasmussen et al., 2011), and reviews emphasize micro-switches and long-range allostery spanning orders of magnitude in time (Miao and McCammon, 2016; Manglik et al., 2015; Venkatakrishnan et al., 2019). By evaluating on A1AR/A2AR, we benchmark whether latent simulators recover metastable states, kinetic pathways, and side-chain signatures relevant to activation.
@@
Recent efforts span coordinate-space trajectory synthesis (e.g., diffusion-based GeoTDM) (Han et al., 2024), SE(3)-equivariant flow matching for coarse-grained rollouts (F3low) (Li et al., 2024), and neural operators for full 3D dynamics beyond next-step prediction (EGNO) (Xu et al., 2024a). Latent and physics-informed approaches accelerate sampling or enforce structure—LAST for adaptive MD (Tian et al., 2023), ConfRover for joint conformation–dynamics learning (Shen et al., 2025), and NeuralMD with symmetric neural ODEs for binding dynamics (Liu et al., 2024)—while comparative studies benchmark diffusion, flow matching, and normalizing flows on MD tasks (John et al., 2025). Complementing these advances, this work isolates the propagator choice by evaluating linear (Koopman), neural, and score-guided dynamics within the same learned latent and fixed decoder, enabling a controlled comparison of long-horizon stability, ensemble fidelity, and functional landscapes. We also considered MD-Gen (Jing et al., 2024), but as it was pre-trained on the Atlas dataset, which does not include alanine-dipeptide, re-training it for a fair comparison was beyond the scope of this work.
@@ 3.4 Score-Guided Langevin Propagator
By evaluating the model at a low, fixed noise level (i.e., a small diffusion timestep $\tau_{\text{noise}}$), we approximate the score of the true data distribution, $s(z)\approx-\frac{\epsilon_{\theta}(z,\tau_{\text{noise}})}{\sigma_{\tau_{\text{noise}}}^{2}}$ (Song et al., 2021; De Bortoli et al., 2021).
@@ 4.2 Equilibrium-ensemble fidelity: NN best on backbone; Langevin best on side-chains
All three models recover the canonical Ramachandran basins. The Autoregressive NN most closely matches basin shape and separation (backbone JSD =0.0056=0.0056), slightly outperforming Koopman ($0.0085$). Langevin under-samples and over-smooths the landscape (backbone JSD =0.029=0.029), consistent with its early rollout failure on this small, rapidly fluctuating system (Fig. 2a). For comparison, an Equivariant Graph Neural Operator (EGNO) baseline (Xu et al., 2024b) yielded a significantly higher backbone JSD of $0.3875$, underscoring the effectiveness of our specialized latent-space propagators.
@@ 4.3 Functional GPCR surfaces: TM-distance free energies capture activation
Table 1: Quantitative summary across systems. Failure time is the first frame with lDDT (vs. the initial frame) <0.65<0.65. JSDs measure divergence from ground-truth dihedral distributions. Average RMSF is in Å. For alanine-dipeptide, we include an EGNO baseline (Xu et al., 2024b) trained for 100 frames. A2AR results are in the SI.

### v2

@@ 1 Introduction
Molecular simulations are indispensable for studying the complex dynamics that govern biological function, yet brute-force approaches struggle to access the slow, functionally relevant motions—such as protein folding, ligand binding, or allosteric switching—due to rugged energy landscapes and the dominance of rare events (Dror et al., 2011; Latorraca et al., 2017). To mitigate this gap—beyond what enhanced sampling can offer when suitable collective variables are hard to specify—a complementary strategy has gained traction: shifting the burden from raw coordinates to learned latent coordinates. In this representation-first view, the simulation problem is recast as a modular encoder–propagator–decoder pipeline: an encoder maps high-dimensional atomic configurations into a continuous, low-dimensional latent space; a propagator evolves the system’s state within this simplified space; and a decoder maps the resulting latent trajectory back to all-atom coordinates (Sidky et al., 2020; Wu et al., 2018).
Progress within this paradigm comes from two complementary directions. The first focuses on learning the underlying physics, employing score-based diffusion, flow matching, and energy-based models to learn generative surrogates—sometimes yielding differentiable force fields—that implicitly define the system’s potential of mean force (Arts et al., 2023; Plainer et al., 2025). The second centers on learning simplified dynamical coordinates, using time-aware autoencoders or Koopman/DMD analysis to discover an intrinsic manifold where the long-term dynamics become stable, predictable, or even approximately linear (Hernández et al., 2018; Mardt et al., 2018; Brunton et al., 2021; Azencot et al., 2020).
@@ 2 Related Work
The core idea of simulating in a low-dimensional space is well established. Molecular Latent Space Simulators (LSS) (Sidky et al., 2020) explicitly factorize the problem into three components: an encoder to find slow collective variables (CVs), a latent propagator to evolve them, and a decoder to generate all-atom structures. Similarly, Deep Generative MSMs (DeepGenMSM) (Wu et al., 2018) pair a latent Markovian transition model with a generative decoder to emit molecular configurations for each state. This blueprint has been realized in various forms, including autoregressive simulators with RNN/LSTM propagators (Vlachas et al., 2021), trajectory-level generators that cast MD as video synthesis (Jing et al., 2024), and invertible models like Boltzmann Generators that learn a direct map to the equilibrium distribution (Noé et al., 2019). We follow this modular design but (i) hold the LD-FPG decoder fixed to control for reconstruction quality and (ii) focus our analysis on the choice of latent propagator.
@@
The quality of a latent simulation hinges on the quality of the latent space itself. While linear methods like time-lagged independent component analysis (TICA) remain strong baselines for identifying slow variables (Pérez-Hernández et al., 2013), deep learning has enabled far more expressive encoders. Time-lagged autoencoders such as the Variational Dynamics Encoder (VDE) learn non-linear representations predictive over a time delay $\Delta t$ (Hernández et al., 2018; Wehmeyer and Noé, 2018). VAMPnets use a variational principle to approximate the leading Koopman eigenfunctions, corresponding to the slowest dynamical processes (Mardt et al., 2018). Similarly, iterative methods like RAVE use VAEs to learn a latent distribution that guides the discovery of an optimal reaction coordinate for enhanced sampling (Ribeiro et al., 2018). To remove hand-crafted features, modern approaches leverage graph neural networks (GNNs) to learn CVs directly from coordinates in a permutation- and symmetry-aware manner (Zou et al., 2025; Ghorbani et al., 2022; Zhang et al., 2024; Duval et al., 2023), often with information-bottleneck objectives to explicitly optimize predictiveness of future states (Wang and Tiwary, 2021).
@@
The Koopman-operator framework provides a route to linearize non-linear dynamics in a learned observable space (Brunton et al., 2021). Extended dynamic mode decomposition (EDMD) (Williams et al., 2015) and DMD (Tu, 2013) showed that a wide class of systems admits accurate linear predictors; subsequent work learns such observables end-to-end with deep encoders (Lusch et al., 2018). This has inspired Koopman autoencoders that enforce a linear evolution rule within the latent space (Tayal et al., 2023; Wu et al., 2025), yielding exceptional long-horizon stability by avoiding the compounding errors of iterated non-linear models, with continued advances in scalable kernels and consistency-enforcing architectures (Azencot et al., 2020).
@@
A complementary path directly learns the non-linear transition function. The Learning Effective Dynamics (LED) framework uses LSTMs to propagate latent variables and capture memory effects (Vlachas et al., 2021); others pair RNNs with physical resampling such as Maximum Caliber to enforce kinetic consistency (Tsai et al., 2022). Continuous-time learners (Neural ODEs) offer flexible parametric flows (Chen et al., 2018), and graph-based simulators exploit locality for rollouts in interacting systems (Sanchez-Gonzalez et al., 2020). To curb instability, physics-preserving architectures (Hamiltonian/Lagrangian networks) enforce conservation laws (Greydanus et al., 2019; Cranmer et al., 2020). Our MLP propagator serves as a streamlined baseline in this family.
@@
Generative models increasingly embed statistical-mechanics structure. Score-based diffusion has been used to learn effective force fields for coarse-grained MD (Arts et al., 2023), linking denoising, Langevin dynamics  (Arts et al., 2023), and Fokker–Planck evolution (Plainer et al., 2025). Recent works (e.g., DiffMD, Score Dynamics) demonstrate larger stable time steps while retaining short-time kinetic signals (Wu and Li, 2023; Hsu et al., 2024). A key theme is consistency between the sampled equilibrium ensemble and the stationary distribution of the learned dynamics (Plainer et al., 2025); stability can also benefit from noise augmentation during training (Arts et al., 2023).
@@
G protein-coupled receptors (GPCRs) are a demanding testbed: their function hinges on slow transitions among metastable states mapped by landmark MD/MSM studies (Dror et al., 2011; Kohlhoff et al., 2014; Latorraca et al., 2017). High-resolution structures (e.g., $\beta_{2}$AR–Gs) provide anchors for validating pathways (Rasmussen et al., 2011), and reviews emphasize micro-switches and long-range allostery spanning orders of magnitude in time (Miao and McCammon, 2016; Manglik et al., 2015; Venkatakrishnan et al., 2016). By evaluating on A1AR/A2AR, we benchmark whether latent simulators recover metastable states and kinetic pathways relevant to activation.
@@
Recent efforts span coordinate-space trajectory synthesis (e.g., diffusion-based GeoTDM) (Han et al., 2024), SE(3)-equivariant flow matching for coarse-grained rollouts (F3low) (Li et al., 2024), and neural operators for full 3D dynamics beyond next-step prediction (EGNO) (Xu et al., 2024). Latent and physics-informed approaches accelerate sampling or enforce structure—LAST for adaptive MD (Tian et al., 2022), ConfRover for joint conformation–dynamics learning (Shen et al., 2025), and NeuralMD with symmetric neural ODEs for binding dynamics (Liu et al., 2024)—while comparative studies benchmark diffusion, flow matching, and normalizing flows on MD tasks (John et al., 2025). Complementing these advances, this work isolates the propagator choice by evaluating linear (Koopman), neural, and score-guided dynamics within the same learned latent and fixed decoder, enabling a controlled comparison of long-horizon stability, ensemble fidelity, and functional landscapes. We also considered MD-Gen (Jing et al., 2024), but as it was pre-trained on the Atlas dataset, which does not include alanine-dipeptide, re-training it for a fair comparison was beyond the scope of this work.
@@ 3.4 Score-Guided Langevin Propagator
By evaluating the model at a low, fixed noise level (i.e., a small diffusion timestep $\tau_{\text{noise}}$), we approximate the score of the true data distribution, $s(z)\approx-\frac{\epsilon_{\theta}(z,\tau_{\text{noise}})}{\sigma_{\tau_{\text{noise}}}^{2}}$ (song2021score; De Bortoli et al., 2021).
@@ 4.2 Equilibrium-ensemble fidelity: NN best on backbone; Langevin best on side-chains
All three models recover the canonical Ramachandran basins. The Autoregressive NN most closely matches basin shape and separation (backbone JSD =0.0056=0.0056), slightly outperforming Koopman ($0.0085$). Langevin under-samples and over-smooths the landscape (backbone JSD =0.029=0.029), consistent with its early rollout failure on this small, rapidly fluctuating system (Fig. 2a). For comparison, an Equivariant Graph Neural Operator (EGNO) baseline (Xu et al., 2024) yielded a significantly higher backbone JSD of $0.3875$, underscoring the effectiveness of our specialized latent-space propagators.
@@ 4.3 Functional GPCR surfaces: TM-distance free energies capture activation
Table 1: Quantitative summary across systems. Failure time is the first frame with lDDT (vs. the initial frame) <0.65<0.65. JSDs measure divergence from ground-truth dihedral distributions. Average RMSF is in Å. For alanine-dipeptide, we include an EGNO baseline (Xu et al., 2024) trained for 100 frames. A2AR results are in the SI.